A 0.275 kg apple falls from a tree to the ground, 3.6 m below. Ignoring air resistance, determine the apple’s kinetic energy, KE, when the apple’s height above the ground is 2.8 m.
1
Expert's answer
2020-04-27T10:14:50-0400

The law of conservation of energy says

"\rm KE_i+PE_i=KE_f+PE_f"

Hence


"{\rm KE_f=KE_i+PE_i-PE_f=0+mgh_i-mgh_f}\\
=0.275\times 9.8\times (3.6-2.8)=2.156\:\rm J"